5. (25 points) Consider an economy described by the production function: Y - F(K, L) K//3, and the depreciation rate is 3 percent. a. Please find the Golden Rule level of capital. b. What is the saving rate that is necessary to reach the Golden Rule level?
Answer #1

a) Golden Rule level of capital is where the marginal productivity of capital equals the depreciation rate.

MPK = Dep. Rate (L is assumed constant)

\partial f/\partial k = 1/3 * k^-2/3 * L^2/3 = 0.03

\Rightarrow k = L/(0.09)^3/2

b) Savings rate needed to achieve Golden Rule level of capital is s

s = Marginal productivity of capital/ Average Producitivity of Capital = 1/3(k)^-2/3*(L)^2/3 / (K^1/3*L^2/3)/K)

= 1/3= 0.33

Savings rate needed is 33.33%,